# Explicit Runge-Kutta Methods with MQ and IMQ-Radial Basis Functions

### v1

@@ Abstract
This article presents a class of explicit Runge-Kutta methods with multiquadric (MQ) and inverse multiquadric (IMQ) radial basis functions (RBFs) to improve the accuracy of time integration for ordinary differential equations. By introducing RBF-based corrections derived from Taylor series expansions and optimally selecting the shape parameter, the method achieves a one-order increase in accuracy without additional stages. Convergence and stability analyses support the theoretical claims, and numerical experiments in MATLAB confirm the predicted performance.
@@ 1 Introduction
where $u(t)$to $C^{\infty}[a,b]$ and $f(t,u)$ is class of $C^{\infty}$ has enormous use in physics, biology, engineering, and control theory. Among the many available techniques, Runge–Kutta (RK) methods are widely used due to their explicit structure, ease of implementation, and high-order accuracy. These methods compute a weighted average of intermediate slope evaluations between $(t_{n},u_{n})$ and $(t_{n+1},u_{n+1})$, and a classical explicit $r$-stage RK method attains order $r$ for $r\leq 4$ [1]. However, for $r\geq 5$, there exists an order barrier, and increasing the number of stages does not necessarily yield higher accuracy if polynomial basis is being used.
To address this limitation, Gu et al. [2] proposed Runge–Kutta methods enhanced with Gaussian radial basis functions (RBFs). Radial basis functions (RBFs) have been integrated into RK methods to enhance their accuracy without increasing the number of stages. This RBF-based method extends the classical approach by embedding a free parameter in the structure of each stage, inspired by the Gaussian RBF Euler method [3]. The idea is to reduce the local truncation error by choosing a free parameter such that the leading error term vanishes, thus improving the order of accuracy by one. The value of the free parameter or the so-called shape parameter demands the local smoothness of the solution.
The shape parameter $\epsilon\in\mathbb{R}\cup i\mathbb{R}$ plays a crucial role in tuning the method both the accuracy and stability of the methods. While there is no universal rule for determining its optimal value, various approaches including stability and accuracy have been used in the literature [2, 3, 4, 5, 6, 7] have even proposed using multiple shape parameters to further enhance accuracy. The use of shape parameters to suppress numerical oscillations or errors is not limited to RBF-based ODE solvers. Similar ideas appear in error inhibiting schemes [8] , Essentially Non-Oscillatory (ENO) and Weighted Essentially Non-Oscillatory (WENO) [9, 10] methods for solving hyperbolic PDEs, where the reconstruction uses adaptive criteria to control local behavior of the solution.
In this work, we adopt the idea of directly eliminating leading truncation error terms through the clever selection of shape parameters. In the literature [4, 5, 6], there are methods in which the optimal value of the shape parameter is used. Here, we analytically determine the shape parameters to cancel terms in the local truncation error expansion up to the desired order. This provides a straightforward and systematic mechanism for constructing high-order RBF-based methods. We extend this framework beyond Gaussian RBFs  [2] and analyze Runge–Kutta methods based on multiquadric (MQ-RBF) [11] and inverse multiquadric (IMQ-RBF) Euler methods [7]. Using the functional form of the RBF-modified Euler methods, we derive the corresponding multistage RK methods and provide local truncation error analysis to establish the convergence of the method. Note that the polynomial-based Runge-Kutta method is a special case of the RBF-based Runge-Kutta method. So, the RBF based Runge-Kutta method shows at least the same order of convergence as that of the classical Runge-Kutta method.
Throughout the paper, $u_{n}$ and $v_{n}$ denote the exact and numerical solutions of the IVP (1.1), respectively. Our analysis focuses on improving the convergence behavior and efficiency of these RBF-RK methods while maintaining explicitness and ease of implementation.
The manuscript is divided into the following sections. Section 2 deals with r-stage classical Runge Kutta method with order r, $r\leq 3$. Section 3 defines the MQ-RBF Runge-Kutta method and the IMQ-RBF Runge-Kutta method and derives the respective expression for the truncation error. In Section 4, we examine the convergence analysis of the MQ and IMQ Runge-Kutta methods. Numerical examples are performed in Section 5, and we demonstrate conclusion and insights in section 6.
2 Explicit Runge-Kutta method
@@
The $r$-stage Runge-Kutta methods with step length $h$ can be written as,
@@
$k_{1}$ $=f(t_{n},v_{n}),$
$k_{2}$ $=f(t_{n}+c_{2}h,v_{n}+ha_{21}k_{1}),$
$k_{3}$ $=f(t_{n}+c_{3}h,v_{n}+h(a_{31}k_{1}+a_{32}k_{2})),$
[…]
$k_{r}$ $=f\left(t_{n}+c_{r}h,v_{n}+h\sum_{j=1}^{r-1}a_{rj}k_{j}\right).$
Explicit r-stage Runge-Kutta methods in Butcher tableau form can be written with Butcher tableau representation as [1]
@@
$c_{r}$ $a_{r1}$ $a_{r2}$dots $a_{r(r-1)}$
$b_{1}$ $b_{2}$dots $b_{r}$
Expanding (2.1) with Taylor’s series and equating the coefficients of h from both sides, we get,
@@
2.1 Two-stage second-order methods
The two-stage Runge-Kutta method can be written as
$v_{n+1}$ $=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}),$ (2.4)
$k_{1}$ $=f(t_{n},v_{n}),$
$k_{2}$ $=f(t_{n}+c_{2}h,v_{n}+ha_{21}k_{1}),$
@@ 2.1 Two-stage second-order methods
along with the conditions (2.2) and (2.3) converts two-stage Runge-Kutta methods to a one-parameter family.
$v^{(1)}$ $=v_{n}+c_{2}hf(t_{n},v_{n}),$
$v_{n+1}$ $=\left(1-\frac{1}{c_{2}}\right)\left(1-\frac{1}{2c_{2}}\right)v_{n}+\frac{1}{c_{2}}\left(1-\frac{1}{2c_{2}}\right)v^{(1)}+\frac{1}{2c_{2}}hf\left(t_{n}+c_{2}h,v^{(1)}\right)$
@@
$\begin{array}[]{c|cc}0&\\ c_{2}&c_{2}\\ \hline\cr&1-\frac{1}{2c_{2}}&\frac{1}{2c_{2}}\\ \end{array}$
when $c_{2}\neq 0.$
2.2 Three-stage third-order methods
Three-stage Runge-Kutta method is of the form
$v_{n+1}$ $=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3}),$ (2.6)
$k_{1}$ $=f(t_{n},v_{n}),$
$k_{2}$ $=f(t_{n}+c_{2}h,v_{n}+ha_{21}k_{1}),$
$k_{3}$ $=f(t_{n}+c_{3}h,v_{n}+h(a_{31}k_{1}+a_{32}k_{2})).$
Butcher tableau representation of the method is given by
$\begin{array}[]{c|ccc}0&\\ c_{2}&a_{21}\\ c_{3}&a_{31}&a_{32}\\ \hline\cr&b_{1}&b_{2}&b_{3}\\ \end{array}$
Another form of three-stage Runge-Kutta method is
$v^{(1)}$ $=v_{n}+c_{2}hf(t_{n},v_{n}),$
$v^{(2)}$ $=\left(1-\frac{a_{31}}{c_{2}}\right)v_{n}+\frac{a_{31}}{c_{2}}v^{(1)}+a_{32}hf\left(t_{n}+c_{2}h,v^{(1)}\right),$
$v_{n+1}$ $=\left(1-\frac{1}{c_{2}}\left(b_{1}-\frac{a_{31}}{a_{32}}b_{2}\right)-\frac{b_{2}}{a_{32}}\right)v_{n}+\frac{1}{c_{2}}\left(b_{1}-\frac{a_{31}}{a_{32}}b_{2}\right)v^{(1)}+\frac{b_{2}}{a_{32}}v^{(2)}+b_{3}hf\left(t_{n}+c_{3}h,v^{(2)}\right).$
@@ 2.2 Three-stage third-order methods
$b_{2}c_{2}+b_{3}c_{3}$ $=\frac{1}{2},$ (2.7)
$b_{2}c_{2}^{2}+b_{3}c_{3}^{2}$ $=\frac{1}{3},$
$a_{32}b_{3}c_{2}$ $=\frac{1}{6},$
along with the conditions (2.2) and (2.3). Based on parameter value the families of three-stage third-order Runge-Kutta methods are given below.
I. $c_{2}\neq 0$, $c_{2}\neq\frac{2}{3}$, $c_{3}\neq 0$, $c_{2}\neq c_{3}$
$\begin{array}[]{c|ccc}0&\\ c_{2}&c_{2}\\ c_{3}&\frac{c_{3}(3c_{3}-3c_{2}^{2}-c_{3})}{c_{2}(2-3c_{2})}&\frac{c_{3}(c_{3}-c_{2})}{c_{2}(2-3c_{2})}\\ \hline\cr&\frac{-3c_{3}+6c_{2}c_{3}+2-3c_{3}}{6c_{2}c_{3}}&\frac{3c_{3}-2}{6c_{2}(c_{3}-c_{2})}&\frac{2-3c_{2}}{6c_{3}(c_{3}-c_{2})}\\ \end{array}$
II. $b_{3}\neq 0$, $c_{2}=c_{3}=\frac{2}{3}$
$\begin{array}[]{c|ccc}0&\\ \frac{2}{3}&\frac{2}{3}\\ \frac{2}{3}&\frac{2}{3}-\frac{1}{4b_{3}}&\frac{1}{4b_{3}}\\ \hline\cr&\frac{1}{4}&\frac{3}{4}-b_{3}&b_{3}\\ \end{array}$
III. $b_{3}\neq 0$, $c_{2}=\frac{2}{3}$,$c_{3}=0$
$\begin{array}[]{c|ccc}0&\\ \frac{2}{3}&\frac{2}{3}\\ 0&-\frac{1}{4b_{3}}&\frac{1}{4b_{3}}\\ \hline\cr&\frac{1}{4}-b_{3}&\frac{3}{4}&b_{3}\\ \end{array}$
@@ 3 Explicit RBF Runge-Kutta Methods
In this section, we present two variants of Runge-Kutta methods with MQ and IMQ-radial basis functions wherein the shape parameter $\epsilon\in\mathbb{R}\cup i\mathbb{R}$. We will further introduce the mathematical expressions for the proposed methods, respective truncation error terms and the optimal value of the shape parameter.
@@ 3.1 Two-Stage Third-Order Methods
(a) MQ
The two-stage MQ-RBF Runge-Kutta method can be written as
$v_{n+1}=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}),$ (3.1)
with,
$k_{1}$ $=f(t_{n},v_{n}),$
$k_{2}$ $=f\left(t_{n}+c_{2}h,(v_{n}+hk_{1}a_{21})\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}\right),$

### v2

# Explicit Runge-Kutta Methods with Multiquadric and Inverse Multiquadric Radial Basis Functions

@@ Abstract
In this article, a family of two- and three-stage explicit multiquadric (MQ) and inverse multiquadric (IMQ) radial basis functions (RBFs) Runge–Kutta methods are introduced for solving ordinary differential equations. These methods are developed by utilizing MQ- and IMQ-RBF Euler methods. The main advantage of these RBF-based methods lies in their ability to achieve a one-order higher accuracy than their classical Runge-Kutta counterparts without increasing the number of stages. This improvement is made possible by incorporating RBF corrections, where the optimal shape parameter is determined through the local truncation error analysis of the proposed schemes. Convergence and stability analyses, including the study of stability regions, are presented to illustrate how these methods compare with standard Runge–Kutta schemes. Numerical experiments on five benchmark problems further confirm predicted accuracy and stability, demonstrating that MQ- and IMQ-based RBF Runge–Kutta methods provide an alternative to conventional low-stage explicit Runge-Kutta schemes.
@@ 1 Introduction
where $u(t)\in C^{\infty}[a,b]$ and $f(t,u)$ is a class of $C^{\infty}$ function has enormous use in physics, biology, engineering, and control theory, etc. Among the many available techniques, Runge–Kutta (RK) methods are widely used due to their explicit structure, ease of implementation, and high-order accuracy. The essential idea behind RK methods is to approximate the solution over a single time step by computing a weighted average of several intermediate slope evaluations of the underlying differential equation. These slope evaluations are taken at carefully chosen points between the current state $(t_{n},u_{n})$ and nect state $(t_{n+1},u_{n+1})$, By combining this information, RK methods construct an improved estimate of the solution trajectory compared to simpler schemes like Euler’s method.
A classical explicit Runge-Kutta method with $r$ stages can, in general, achieve accuracy of order $r$ when $r\leq 4$ [1]. However, for $r\geq 5$, there arises a well-known order barrier: the relationship between the number of stages and the achievable order of accuracy becomes nonlinear. In particular, beyond fourth order, simply increasing the number of stages within a polynomial-based framework does not automatically yield a corresponding increase in accuracy. The limitation motivates the exploration of alternative formulations and basis functions in the design of higher-order integrators. For implicit Runge-Kutta methods, a $r$- stage scheme can reach order $2r,$ that goes beyond the order limitations of explicit methods, though at the expense of solving nonlinear systems at each step. Over time, various specialized Runge-Kutta schemes have been developed for different applications: exponential-fitted schemes for oscillatory problems [2, 3, 4], symplectic schemes for Hamiltonian systems [5, 6, 7], exponential schemes for stiff systems [8, 9, 10], total variation diminishing and strong stability preserving schemes for hyperbolic conservation laws [11, 12, 13, 14], and implicit–explicit schemes for problems involving both stiff and non-stiff components [15, 16, 17]. These variants demonstrate the flexibility and broad applicability of the Runge–Kutta framework.
The primary objective of this work is to explore the role of radial basis functions (RBFs) in the construction of accurate time discretization schemes for the numerical solution of differential equations. To set the stage, we first review relevant contributions from the existing literature. In [18, 19], adaptive RBF-based solvers were introduced for solving initial value problems (IVPs). These solvers reformulated classical time-stepping methods—such as Euler’s method, the midpoint method, and the Adams family of methods—by replacing the traditional polynomial bases with radial basis functions, specifically Multi-Quadratic (MQ) and Gaussian RBFs. The distinguishing feature of such adaptive RBF solvers lies in the presence of a free shape parameter embedded in the basis functions. This parameter is not fixed but can be adaptively tuned according to the local smoothness of the solution, thereby enhancing accuracy and stability. A rigorous strategy for selecting this parameter involves using derivative information to eliminate the leading terms in the local truncation error expansion, effectively improving the order of accuracy of the method. Building upon this foundation, subsequent works such as [20, 21] extended these ideas further by designing time discretization schemes based on inverse quadratic and inverse multi-quadratic RBFs, thereby broadening the applicability of RBF-based solvers across a wider range of problems.
The shape parameter $\epsilon$ is central to balancing accuracy and stability in RBF-based methods. Since no universal rule exists for its optimal choice, various strategies based on stability and accuracy have been explored [22, 23, 24], including the use of multiple shape parameters to further improve accuracy. The use of shape parameters to suppress numerical oscillations or errors is not limited to RBF-based ODE solvers. Similar ideas appear in error inhibiting schemes [25], Essentially Non-Oscillatory (ENO) and Weighted Essentially Non-Oscillatory (WENO) [26, 27] methods for solving hyperbolic PDEs, where the reconstruction uses adaptive criteria to control local behavior of the solution. To advance the construction of higher-order time discretization schemes, Gu et al. [28] introduced Runge–Kutta (RK) methods augmented with Gaussian radial basis functions (RBFs). In this framework, RBFs are incorporated into the stage equations of RK methods to enhance their accuracy without increasing the number of stages. The resulting RBF-embedded RK schemes generalize the classical formulation by introducing a free shape parameter at each stage, motivated by the Gaussian RBF modification of Euler’s method [19]. The central idea is to systematically reduce the local truncation error by selecting the shape parameter such that the dominant error term in the expansion is canceled, thereby raising the effective order of accuracy by one. The determination of this parameter depends on the local smoothness of the underlying solution, which guides its adaptive adjustment during the integration process.
In this work, we propose a family of two- and three-stage explicit multiquadric (MQ) and inverse multiquadric (IMQ) radial basis function RK methods for solving ordinary differential equations. These methods are developed by extending MQ- and IMQ-based RBF Euler methods, and are designed to achieve one order higher accuracy than classical RK methods without increasing the number of stages. This enhancement is made possible by incorporating RBF corrections, where the shape parameters are analytically selected to eliminate leading terms in the local truncation error expansion, offering a systematic approach to constructing high-order RBF-based schemes. Unlike prior works [22, 23, 24] that rely on empirical or numerically optimized shape parameters, this approach determines optimal values through rigorous local error analysis. The proposed framework generalizes beyond Gaussian RBFs, covering both MQ and IMQ variants, and includes detailed convergence and stability analyses such as the study of stability regions to assess the performance relative to standard RK schemes. Numerical experiments on five benchmark problems confirm the predicted accuracy and stability, demonstrating that these RBF-based methods not only match but often exceed the performance of conventional low-stage explicit RK methods. Notably, classical polynomial-based RK methods emerge as special cases of this broader RBF-based formulation, ensuring that the new methods maintain at least the same order of convergence while offering enhanced flexibility and accuracy.
The manuscript is organized as follows; Section 2 reviews classical $r$ stage Runge–Kutta methods of order $r\leq 3$ laying the groundwork for comparison with the proposed methods. Section 3 introduces the MQ-RBF and IMQ-RBF Runge–Kutta methods, deriving their formulations from the corresponding RBF-modified Euler methods. We perform local truncation error analysis and establish order conditions to ensure that an $r$ stage method achieves $(r+1)$-th order accuracy through optimal shape parameter selection. Section 4 presents the convergence analysis of the proposed methods, demonstrating their theoretical consistency and accuracy. Section 5 provides numerical experiments on benchmark problems to validate the improved performance and stability of the RBF-based methods. Finally, some conclusions and future works are provided in Section 6.
2 Explicit Runge-Kutta Methods
@@
We denote $u_{n}$ and $v_{n}$ as the exact and numerical solutions of the IVP (1.1), respectively. The $r$-stage Runge-Kutta methods [1, 29] with step length $h$ for (1.1) can be written as,
@@
$\begin{aligned} k_{1}&=f(t_{n},v_{n}),\\ k_{2}&=f(t_{n}+c_{2}h,v_{n}+ha_{21}k_{1}),\\ k_{3}&=f(t_{n}+c_{3}h,v_{n}+h(a_{31}k_{1}+a_{32}k_{2})),\\ &\ \vdots\\ k_{r}&=f\left(t_{n}+c_{r}h,v_{n}+h\sum_{j=1}^{r-1}a_{rj}k_{j}\right).\end{aligned}$
are slopes at intermediate points $t_{n}+c_{r}h$ between $t_{n}$ to $t_{n+1},$ $a_{rj}$ for $1\leq j<i\leq r$ are coefficients, $b_{i}$ for $1\leq i\leq r,$ and $c_{r}$ for $1\leq i\leq r.$ The matrix $a_{rj}]$ is called the Runge-Kutta matrix, $b_{i}$ are known as weights and $c_{r}$ are as nodes. Such r-stage explicit Runge-Kutta method can be written in Butcher tableau form [30] as follows
@@
$c_{r}$ $a_{r1}$ $a_{r2}$ $\dots$ $a_{r(r-1)}$
$b_{1}$ $b_{2}$ $\dots$ $b_{r}$
Expanding (2.1) with Taylor’s series and equating the same order coefficients of h for both sides, we get consistency conditions
@@
and to get required higher-order accuracy for this work, we assume that it satisfies the following condition;
@@
Two-stage second-order methods: The two-stage Runge-Kutta methods can be written as
$\begin{cases}\begin{aligned} v_{n+1}&=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}),\\ k_{1}&=f(t_{n},v_{n}),\\ k_{2}&=f(t_{n}+c_{2}h,v_{n}+ha_{21}k_{1}),\end{aligned}\end{cases}$ (2.4)
@@
along with the conditions (2.2) ($b_{1}+b_{2}=1$) and (2.3) ($a_{21}=c_{2}$) converts two-stage Runge-Kutta methods to a one-parameter family of schemes with $c_{2}\neq 0$ as
$\begin{cases}\begin{aligned} v^{(1)}&=v_{n}+c_{2}hf(t_{n},v_{n}),\\ v_{n+1}&=\left(1-\frac{1}{c_{2}}\left(1-\frac{1}{2c_{2}}\right)\right)v_{n}+\frac{1}{c_{2}}\left(1-\frac{1}{2c_{2}}\right)v^{(1)}+\frac{1}{2c_{2}}hf\left(t_{n}+c_{2}h,v^{(1)}\right).\end{aligned}\end{cases}$
@@
[…]
Three-stage third-order methods: The three-stage Runge-Kutta methods can be expressed in the form
$\begin{cases}\begin{aligned} v_{n+1}&=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}+b_{3}k_{3}),\\ k_{1}&=f(t_{n},v_{n}),\\ k_{2}&=f(t_{n}+c_{2}h,v_{n}+ha_{21}k_{1}),\\ k_{3}&=f(t_{n}+c_{3}h,v_{n}+h(a_{31}k_{1}+a_{32}k_{2})).\end{aligned}\end{cases}$ (2.6)
Butcher tableau representation of this method is given by
[…]
Alternative form of three-stage Runge-Kutta method is
$\begin{cases}\begin{aligned} v^{(1)}&=v_{n}+c_{2}hf(t_{n},v_{n}),\\ v^{(2)}&=\left(1-\frac{a_{31}}{c_{2}}\right)v_{n}+\frac{a_{31}}{c_{2}}v^{(1)}+a_{32}hf\left(t_{n}+c_{2}h,v^{(1)}\right),\\ v_{n+1}&=\left(1-\frac{1}{c_{2}}\left(b_{1}-\frac{a_{31}}{a_{32}}b_{2}\right)-\frac{b_{2}}{a_{32}}\right)v_{n}+\frac{1}{c_{2}}\left(b_{1}-\frac{a_{31}}{a_{32}}b_{2}\right)v^{(1)}\\ &\qquad+\frac{b_{2}}{a_{32}}v^{(2)}+b_{3}hf\left(t_{n}+c_{3}h,v^{(2)}\right).\end{aligned}\end{cases}$
@@
$b_{2}c_{2}+b_{3}c_{3}=\frac{1}{2},\,\,\,b_{2}c_{2}^{2}+b_{3}c_{3}^{2}=\frac{1}{3},\,\,a_{32}b_{3}c_{2}=\frac{1}{6},$ (2.7)
along with the conditions (2.2) and (2.3). The possible families of three-stage third-order Runge-Kutta methods [30] are given below.
I. $c_{2}\neq 0$,  $c_{2}\neq\frac{2}{3}$,  $c_{3}\neq 0$,   $c_{2}\neq c_{3}$.
[…]
II. $b_{3}\neq 0$,  $c_{2}=c_{3}=\frac{2}{3}$.
[…]
III. $b_{3}\neq 0$,  $c_{2}=\frac{2}{3}$,  $c_{3}=0$.
[…]
@@ 3 Explicit RBF Runge-Kutta Methods
In this section, we present two variants of Runge-Kutta methods with multi-quadratic (MQ) and inverse multi-quadratic (IMQ) radial basis functions based on respective Euler methods.
[…]
where the shape parameter $\epsilon\in\mathbb{R}\cup i\mathbb{R}$. Under special conditions, the square root involved in the above methods may be difficult to evaluate, thus the authors presented remedies [18, 20]. Such methods read as
[…]
Now, we focus on developing mathematical expressions for the MQ and IMQ-RBF based Runge-Kutta schemes with the building block of Euler methods in (3.1) rather than in (3.2). We further analyze truncation errors and derive the order of the method according to the consistency conditions.
@@ 3.1 Two-Stage Third-Order Methods
In this subsection, we develop two-stage third-order MQ and IMQ-RBF based Runge-Kutta schemes and perform the turncation error analysis.
(a) MQ-RBF RK scheme: The two-stage MQ-RBF Runge-Kutta method can be written as
$\begin{cases}\begin{aligned} v_{n+1}&=v_{n}+h(b_{1}k_{1}+b_{2}k_{2}),\\ k_{1}&=f(t_{n},v_{n}),\\ k_{2}&=f\left(t_{n}+c_{2}h,(v_{n}+hk_{1}a_{21})\sqrt{1+\epsilon_{n}^{2}a_{21}^{2}h^{2}}\right),\end{aligned}\end{cases}$ (3.3)